The sum of the first n positive odd integers is n2.

Establish the formula for the sum of the first n positive even integers and use proof by mathematical induction to prove its correctness.


1
Expert's answer
2021-10-15T10:12:45-0400

The sum of the first "n" positive even integers is "n(n+1)."

Let "P(n)" be the proposition that the sum of the first n positive even integers is "n(n+1)."

Basic Step: "P(1)" is true because "2=1(1+1)."

Inductive Step: For the inductive hypothesis we assume that "P(k)" holds for an arbitrary

positive integer "k." That is, we assume that


"2+4+...+2k=k(k+1)"

Under this assumption, it must be shown that "P(k + 1)" is true, namely, that


"2+4+...+2k+2(k+1)=(k+1)(k+1+1)"

is also true.

When we add "2(k + 1)" to both sides of the equation in "P(k)," we obtain


"2+4+...+2k+2(k+1)=k(k+1)+2(k+1)"

"=(k+1)(k+2)"

This last equation shows that "P(k + 1)" is true under the assumption that "P(k)" is true. This completes the inductive step.

We have completed the basis step and the inductive step, so by mathematical induction we know that "P(n)" is true for all positive integers n. That is, we have proved that


"2+4+...+2n=n(n+1)"

for all positive integers n.

The sum of the first "n" positive even integers is "n(n+1)."